Problem 7-5B Determine depreciation under three methods (LO7-4)

[The following information applies to the questions displayed below.]

Cheetah Copy purchased a new copy machine. The new machine cost $134,000 including installation. The company estimates the equipment will have a residual value of $33,500. Cheetah Copy also estimates it will use the machine for four years or about 8,000 total hours. Actual use per year was as follows:

Required information Problem 7-5B Determine depreciation under three methods (LO7-4) [The following information applies to thProblem 7-5B Part 1 Required: 1. Prepare a depreciation schedule for four years using the straight-line method. (Do not roundcompany estimates the equipment will have a residual value of $33,500. Cheetah Copy also estimates it will use the machine foProblem 7-5B Part 2 2. Prepare a depreciation schedule for four years using the double-declining-balance method. (Hint: The aProblem 7-5B Part 3 3. Prepare a depreciation schedule for four years using the activity-based method. (Round your Depreciat
Answer #1
Solution a
Cost of assets $ 134,000.00
Salvage value $   33,500.00
Depreciable value $ 100,500.00
SLM depreciation= Depreciable value/useful life
SLM depreciation= 100500/4 $                               25,125.00
Year Depreciation Accumulated depreciation Book Value
1 $         25,125 $                                     25,125 $      108,875
2 $         25,125 $                                     50,250 $         83,750
3 $         25,125 $                                     75,375 $         58,625
4 $         25,125 $                                  100,500 $         33,500
Solution b
Useful life                       4
Annual SLM depreciation 1/4 25%
Double declining depreciation 50.00%
Year Beginning book value Depreciation Depreciation Accumulated depreciation Yearend Book Value
A B C= A*B D E
1 $      134,000 50% $         67,000 $   67,000 $   67,000
2 $         67,000 50% $         33,500 $ 100,500 $   33,500
3 $         33,500 50% $                 -   $ 100,500 $   33,500
4 $         33,500 50% $                 -   $ 100,500 $   33,500
Since book value has already reached salvage value hence no further depreciation will be charged in years 3 and 4.
Solution C
Calculation of depreciation rates
Year Activity Percentage
Annual activity/Total activity
1               2,800 28.00%
2               2,000 20.00%
3               2,000 20.00%
4               3,200 32.00%
            10,000
Year Depreciable value Depreciation Depreciation Accumulated depreciation Yearend Book Value
A B C= A*B D E
1 $      100,500 28% $         28,140 $   28,140 $ 105,860
2 $      100,500 20% $         20,100 $   48,240 $   85,760
3 $      100,500 20% $         20,100 $   68,340 $   65,660
4 $      100,500 32% $         32,160 $ 100,500 $   33,500